MsBrains (ms-brains):

PLEASE HELP ME BY CHECKING IF MY ANSWER IS CORRECT, AND HELPING ME UNDERSTAND!! I WILL MEDAL!

MsBrains (ms-brains):

I don't know this one, Can some explain how to solve?

OpenStudy (agl202):

2x-y=0 2(1) - (2) = 0 2 - 2 = 0=0 So it eventually be ?

MsBrains (ms-brains):

0? o.O

OpenStudy (agl202):

U will be replacing the x and y with (1,2) Like I showed ^

MsBrains (ms-brains):

Oh, so it would be B! (1,2)

OpenStudy (agl202):

Ye :) Hopefully u understood.

MsBrains (ms-brains):

I did! Can you help me with some more?

OpenStudy (agl202):

Sure

MsBrains (ms-brains):

I think it is A

OpenStudy (agl202):

Do u understand how to read graphs?

MsBrains (ms-brains):

Yes, and I think the answer is A..?

OpenStudy (agl202):

Ur right.

MsBrains (ms-brains):

D?

OpenStudy (agl202):

No, not D.

MsBrains (ms-brains):

What it is it then? I thought since it was passing through -4

MsBrains (ms-brains):

Oh! so it's B?

OpenStudy (agl202):

You got it :)

MsBrains (ms-brains):

How do I solve this one..?

OpenStudy (agl202):

You will use this for your coordinates: https://encrypted-tbn0.gstatic.com/images?q=tbn:ANd9GcQN0h51BA_0Mcu3eA4GxHtfjcQk6oEHbFWSkhTRZVTD5S6RfSdamQ You will get: 5-3/2-1 =slope 2/1 =2

OpenStudy (agl202):

Wait. My bad.

OpenStudy (agl202):

Sorry I was lagging. I forgot the negative^

OpenStudy (agl202):

So u will be getting -4/5.

OpenStudy (agl202):

what do u mean by 2?

MsBrains (ms-brains):

Nvm. Okay got it! I have 9 questions. If you are free

OpenStudy (agl202):

What Q r we on now?

OpenStudy (agl202):

It better not be any tests cuz that's against the CoC. :/

MsBrains (ms-brains):

Its not a test! It's for studying. :) we are on 5/14

OpenStudy (agl202):

o.O K

MsBrains (ms-brains):

I think it's C

MsBrains (ms-brains):

I mean D. I think it is D. :)

OpenStudy (agl202):

Wait.

OpenStudy (agl202):

Actually it is D.

MsBrains (ms-brains):

Yay! Ok. Next one

OpenStudy (agl202):

Slope intercept form is y=mx+b m is slope, b is y-intercept 4x - 9y = -36 4x - 9y + 9y = -36 + 9y 4x = -36 + 9y 4x + 36 = -36 + 9y + 36 4x + 36 = 9y 9y = 4x + 36 9y/9 = 4x/9 + 36/9 y = (4/9)x + 4

MsBrains (ms-brains):

C.

OpenStudy (agl202):

Yep

OpenStudy (agl202):

What do u think?

MsBrains (ms-brains):

I think it's B?

OpenStudy (agl202):

Me 2

MsBrains (ms-brains):

Are you sure? How can we check it's for certain..?

OpenStudy (agl202):

K

OpenStudy (agl202):

The lines will be the same, actually.

MsBrains (ms-brains):

So it was C?

OpenStudy (agl202):

Ye

OpenStudy (agl202):

I am thinking B. for those negatives

MsBrains (ms-brains):

Are you sure?

OpenStudy (agl202):

Lemme check

OpenStudy (agl202):

B it is

OpenStudy (agl202):

Where r we at now?

MsBrains (ms-brains):

Question 9. :)

MsBrains (ms-brains):

So what do you think? ^

MsBrains (ms-brains):

@Agl202 ?? you there

OpenStudy (agl202):

I'm sorry but i gtg :/

MsBrains (ms-brains):

Oh! It's okay! See ya :)